OpenStudy (anonymous):

Jamie invested 800 at 8.2% annual interest compounded continuously. What is the final amount of her investment after 20 years?

OpenStudy (anonymous):

compounded continuosly? Are you sure it's not annually?

OpenStudy (anonymous):

well its simple formula =800*0.082*20

OpenStudy (anonymous):

It is compounded continuously.

OpenStudy (anonymous):

I=P r T ........ P priceinitial , r rate in and T time .. r is given in percentage .. divivded it by 100

OpenStudy (anonymous):

I failed. I got way too low of a number.

OpenStudy (anonymous):

so =800*0.082*20 =1312

OpenStudy (anonymous):

Yay but the options are all bigger than that.

OpenStudy (anonymous):

A. 3869.32 B. 4124.14 C. 4356.63 D. 4409.87

OpenStudy (anonymous):

http://math.about.com/od/businessmath/ss/Interest.htm you should take help from here man

OpenStudy (anonymous):

A = P(1 + r)n

OpenStudy (neer2890):

\[C.I.=P(1+ \frac{ r }{ 100 })^{T}\]

OpenStudy (neer2890):

where p= principle r= rate of interest per annum and T= time in years

OpenStudy (anonymous):

is it .082 or 8.2?

OpenStudy (anonymous):

800*(1.082)^20 3869

OpenStudy (anonymous):

Thanks

OpenStudy (neer2890):

so it is 3869.32

OpenStudy (anonymous):

Thanks to you both. You have saved me for finals tomorrow. :)

OpenStudy (neer2890):

you're welcome..:)

OpenStudy (anonymous):

What about if 3,000 at 8.2% annual interest is compounded montly? How much would the investment be after 15 years?

OpenStudy (neer2890):

\[C.I.= P(1+\frac{ \frac{ r }{ m } }{ 100 })^{mT}\] where P,r and T are same as in first question and m is no. of time periods in one year.

OpenStudy (neer2890):

for yearly m=1 for half yearly m=2 for quarterly, m=3 for monthly, m=12 for 2 yearly m=1/2 for t yearly m=1/t

OpenStudy (anonymous):

So what would that end up as?